# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 4       $k\leftarrow k+1$;
Remark 2.13 (Without Total Boundedness).
For the upper bound, the algorithm may require solving an optimization over infinitely many functions, but our results still hold. Lemma 2.5 and Lemma 2.6 do not rely on any packing constructions but simply a concentration result. Lemma 2.7 is a statement about any packing set, not necessarily a maximal or infinite one. Monotonicity of the local metric entropy is unaffected. The arguments within the proof of Theorem 2.11 and Theorem 2.12 also hold either trivially (e.g., that the probability is bounded by infinity) or a key hypothesis such as (2.6) implies the local metric entropy is finite.
The similar proofs in the following unbounded and adaptive settings will also be unaffected.
3 Extension to the Unbounded Case
Let us now assume $\mathcal{F}$ is not bounded in sup-norm (i.e., for all $r>0$, there is an $f\in\mathcal{F}$ and $x\in\mathcal{X}$ such that $|f(x)|>r$). But recall that we assume the diameter $d:=\mathrm{diam}(\mathcal{F}):=\sup\{\|f-g\|_{L_{2}(\mathbb{P}_{X})}:f,g\in\mathcal{F}\}$ is bounded.
However, we require an additional assumption similar to mendelson_2017_local_versus_global, namely, that there is some $\alpha>0$ such that for each $p\geq 2$ and $f,g\in\mathcal{F}$ we have
$\|f-g\|_{L_{p}(\mathbb{P}_{X})}\leq\alpha\sqrt{p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}.$ (3.1)
Here we define for any $p\geq 2$ the $L_{p}(\mathbb{P}_{X})$-distance between $f,g\in\mathcal{F}$ by
$\|f-g\|_{L_{p}(\mathbb{P}_{X})}=\left(\int|f(x)-g(x)|^{p}\mathrm{d}\mathbb{P}_{X}(x)\right)^{1/p}.$
We omit the assumption about boundedness of the induced centered canonical Gaussian process $\{G_{f}:f\in\mathcal{F}\}$ used in mendelson_2017_local_versus_global. Rather, we assume boundedness of the diameter. Moreover, our approach is able to match our lower and upper bounds.
Our lower bounds from Section 2 all carry through with no change. For the upper bound, we will need slightly different concentration inequalities, but several of the proofs will carry over with minimal changes.
Define the subexponential norm $\|\cdot\|_{\psi_{1}}$ of a random variable $Z$ by $\|Z\|_{\psi_{1}}=\sup_{p\geq 1}p^{-1}\cdot\left(\mathbb{E}|Z|^{p}\right)^{1/p}.$ Then using Minkowski’s inequality and (3.1), we have
$\left|(f(X)-g(X))^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right|_{\psi_{1}}$ $=\sup_{p\geq 1}p^{-1}\sqrt[p]{\mathbb{E}|(f(X)-g(X))^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}|^{p}}$
$\leq\sup_{p\geq 1}p^{-1}\left[\sqrt[p]{\mathbb{E}|f(X)-g(X)|^{2p}}+\sqrt[p]{\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2p}}\right]$
$=\sup_{p\geq 1}p^{-1}\left[\left(\sqrt[2p]{\mathbb{E}|f(X)-g(X)|^{2p}}\right)^{2}+\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right]$
$=\sup_{p\geq 1}p^{-1}\left[\|f-g\|_{L_{2p}(\mathbb{P}_{X})}^{2}+\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right]$
$\leq\sup_{p\geq 1}p^{-1}\left(2\alpha^{2}p+1\right)\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$\leq(2\alpha^{2}+1)\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$\leq(2\alpha^{2}+1)d\|f-g\|_{L_{2}(\mathbb{P}_{X})}$
$\leq(2\alpha^{2}+1)d^{2}.$ (3.2)
We now use this result along with concentration inequalities for subexponential random variables from vershynin_2012 to obtain the following analogue of Lemma 2.5. As before, the result holds when the noise is sub-Gaussian.
Lemma 3.1.
Suppose $\mathcal{F}$ satisfies (3.1), and $f,g\in\mathcal{F}$ satisfy $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some sufficiently large $C>0$. Suppose further that $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. Then $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$ both simultaneously hold with probability at least
$1-2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right),$
where $b>0$ is some universal constant.
{toappendix}
Proof of Lemma 3.1.
Consider the random variables $Z_{i}=(f(X_{i})-g(X_{i}))^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$, which are centered and sub-exponential (vershynin_2012, Definition 5.13 or property (5.15)) with $\|Z_{i}\|_{\psi_{1}}\leq(2\alpha^{2}+1)d^{2}$ for each $i$ by (3.2). Set $K:=\max_{1\leq i\leq n}\|Z_{i}\|_{\psi_{1}}$. By Proposition 5.16 of vershynin_2012, we have for some absolute constant $b>0$ that
$\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}<-t\right)$ $\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
Then setting $t=C^{-2}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$,
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$ $=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot C^{2}\delta^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}<-t\right)$
$\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
Observe that since $K\leq(2\alpha^{2}+1)d\|f-g\|_{L_{2}(\mathbb{P}_{X})}$ (see (3.2)),
$\frac{t^{2}}{nK^{2}}$ $=\frac{C^{-4}n^{2}\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{4}}{nK^{2}}\geq\frac{C^{-4}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{4}}{(2\alpha^{2}+1)^{2}d^{2}\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}=\frac{C^{-4}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}{(2\alpha^{2}+1)^{2}d^{2}}$
$\geq\frac{C^{2}\delta^{2}}{C^{4}(2\alpha^{2}+1)^{2}d^{2}}=\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}.$
Similarly, since $K\leq(2\alpha^{2}+1)d^{2}$,
$\frac{t}{K}$ $=\frac{C^{-2}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}{K}\geq\frac{C^{-2}\cdot C^{2}\delta^{2}}{(2\alpha^{2}+1)d^{2}}=\frac{\delta^{2}}{(2\alpha^{2}+1)d^{2}}\geq\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}},$
so $\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\geq\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}$. Therefore,
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$ $\leq\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Next, we consider the random variables $Z_{i}=(f(X_{i})-\bar{f}(X_{i}))^{2}-\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$, which are again centered and sub-exponential with $\|Z_{i}\|_{\psi_{1}}\leq(2\alpha^{2}+1)d^{2}$ for each $i$. Let $K$ be defined as before. By Proposition 5.16 of vershynin_2012, we have for the same absolute constant $b>0$ that
$\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}>t\right)$ $\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
We set $t=\delta^{2}$ and obtain
$=\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}>t\right)$
$\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
Since $\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\delta/\sqrt{n}$, we have
$K\leq(2\alpha^{2}+1)d\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{(2\alpha^{2}+1)d\delta}{\sqrt{n}}.$
Thus,
$\frac{t^{2}}{nK^{2}}$ $=\frac{\delta^{4}}{nK^{2}}\geq\frac{n\delta^{4}}{n(2\alpha^{2}+1)^{2}d^{2}\delta^{2}}=\frac{\delta^{2}}{(2\alpha^{2}+1)^{2}d^{2}}.$
Further, since $K\leq(2\alpha^{2}+1)d^{2}$,
$\frac{t}{K}$ $=\frac{\delta^{2}}{K}\geq\frac{\delta^{2}}{(2\alpha^{2}+1)d^{2}}\geq\frac{\delta^{2}}{(2\alpha^{2}+1)^{2}d^{2}}.$
Hence $\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\geq\frac{\delta^{2}}{(2\alpha^{2}+1)^{2}d^{2}}\geq\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}$. Therefore,
$\mathbb{P}_{\bar{f}}\left(n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}>2\delta^{2}\right)$ $\leq\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Combining these results with a union bound, we conclude that with probability at least
$1-2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right),$
we have $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$. ∎
We then obtain analogues of Lemmas 2.6 and 2.7.
Lemma 3.2.
Suppose $\mathcal{F}$ satisfies (3.1), and suppose we are testing $H_{0}:\bar{f}=f$ vs $H_{1}:\bar{f}=g$ for $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some $C>3$. Then for some universal constant $b>0$, the test $\psi(\vec{Y})=\mathbbm{1}\left(\|\vec{Y}-f(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-g(\vec{X})\|_{2}^{2}\right)$ satisfies
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-\widetilde{L}(C,\sigma)\cdot\delta^{2}\right),$
where
$\widetilde{L}(C,\sigma)=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{b}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right\}.$ (3.3)
{toappendix}
Proof of Lemma 3.2.
Many of the steps are identical to those from the proof of Lemma 2.6, so we omit some details. The main difference is in the specific probability bounds we obtain. Let $\eta(\vec{X})=\bar{f}(\vec{X})-f(\vec{X})$. Then as detailed in Lemma 2.6, we derive
$\quad\quad\quad+2\xi^{T}(g(\vec{X})-f(\vec{X})).$
and variance $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$.
We now derive estimates on $\|\eta(\vec{X})\|_{2}$ and $\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ that hold with high probability. Suppose $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. From Lemma 3.1, this condition along with our assumption $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ implies that with probability at least $1-2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right)$, we have
$n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2},\quad n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}.$ (3.4)
Let $E$ denote the event that both inequalities of (3.4) hold, so
$\mathbb{P}_{\bar{f}}(E^{c})\leq 2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Note that the inequalities in (3.4) are the same two inequalities as in the proof of Lemma 2.6.
Hence following the details in Lemma 2.6, we obtain that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
Since $\mathbb{P}_{\bar{f}}(\psi=1|\vec{X},E)$ has an upper bound not depending on $X$, it follows (by taking an expectation) that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
Thus,
$\mathbb{P}_{\bar{f}}(\psi=1)$ $=\mathbb{P}_{\bar{f}}(\psi=1|E)\cdot\mathbb{P}_{\bar{f}}(E)+\mathbb{P}_{\bar{f}}(\psi=1|E^{c})\cdot\mathbb{P}_{\bar{f}}(E^{c})$
$\leq\mathbb{P}_{\bar{f}}(\psi=1|E)+\mathbb{P}_{\bar{f}}(E^{c})$
$\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
If we instead assumed $n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$, we would arrive on an identical bound for $\mathbb{P}_{\bar{f}}(\psi=0)$. Thus, we conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq$
$\quad\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
We define
$\widetilde{L}(C,\sigma)=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{b}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right\},$
to conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-\widetilde{L}(C,\sigma)\cdot\delta^{2}\right).$
Lemma 3.3.
Suppose $\mathcal{F}$ satisfies (3.1). Let $f_{1},\dots,f_{M}$ be a $\delta$-packing of $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ and suppose $\bar{f}\in\mathcal{F}^{\prime}$. Let $i^{\ast}=\mathop{\mathrm{argmin}}_{i\in[M]}\sum_{j=1}^{n}(Y_{j}-f_{i}(X_{j}))^{2}.$ Then
$\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1)\right)$ $\leq 3M\exp\left(-\widetilde{L}(C,\sigma)\cdot\delta^{2}\right),$
where $\widetilde{L}(C,\sigma)$ is defined as in (3.3).
{toappendix}
Proof of Lemma 3.3.
The proof is completely identical to that of Lemma 2.7, except in the step (2.5), we instead use Lemma 3.2. This replaces the $L(C,\sigma,B_{\mathcal{F}})$ from the bounded case with $\widetilde{L}(C,\sigma)$ where $\widetilde{L}(C,\sigma)$ is defined in (3.3).
Remark 3.4.
Lemmas 3.2 and 3.3 can be generalized to our sub-Gaussian noise setting where each $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$. We supply the details in the appendix.
{toappendix}
Proof of Remark 3.4.
As in Lemma 3.2, we show that
$\quad\quad\quad+2\xi^{T}(g(\vec{X})-f(\vec{X})).$
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2},$
and has variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. The proof of Lemma 3.2 then defines the event $E$ as when the inequalities in (3.4) hold, and $E$ satisfies
$\mathbb{P}_{\bar{f}}(E^{c})\leq 2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
$\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$
and variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. A sub-Gaussian random variable $Z$ with mean $\rho<0$ and variance proxy $\tau^{2}$ can be shown to satisfy $P(Z>0)\leq\exp(-\rho^{2}/2\tau^{2})$ using arguments from boucheron2013. Applying this result, as in the previous lemma, we conclude that conditional on $\vec{X}$ and the event $E$, $\psi=1$ holds with probability bounded by
$\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
The rest of the proof follows identically to Lemma 3.2. Then we use this generalized lemma to obtain Lemma 3.3 since normality is not used. ∎
Thus we obtain similar algorithmic guarantees in the following Theorem. By Remark 3.4, the proof holds for sub-Gaussian noise as well.
Theorem 3.5.
Suppose $\mathcal{F}$ satisfies (3.1). Let $J^{\ast}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}=\frac{d\sqrt{\widetilde{L}(c/2-1,\sigma)}}{2^{J-2}c}$ satisfying
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{\varepsilon_{J}\cdot c}{\sqrt{\widetilde{L}(c/2-1,\sigma)}},c\right)\vee\log 2.$ (3.5)
and $J^{\ast}=1$ if this condition is never satisfied. Let $f^{\ast}(\vec{X},\vec{Y})$ be the output of $J^{\ast}$ steps of Algorithm 1. Then $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}\varepsilon_{J^{\ast}}^{2}$ for some constant $\overline{C}$ depending only on universal constants and $\sigma$.We defined $\widetilde{L}(C,\sigma)$ in (3.3), where $C$ is some sufficiently large universal constant and $c:=2(C+1)$ is the constant we use in the definition of local metric entropy.
{toappendix}
@@ 3 Extension to the Unbounded Case
Suppose $J>1$ steps of the algorithm are performed and consider $2\leq j\leq J$. Let $\Upsilon_{j+1}$ be the function added to $\Upsilon$ in step $j$. During step $j-1$, we start with some $\Upsilon_{j-1}$ and then form a $\frac{d}{2^{j-1}(C+1)}$-packing $M_{j-1}$ of the set $B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$. We then set $\Upsilon_{j}=\mathop{\mathrm{argmin}}_{f\in M_{j-1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$. Applying Lemma 3.3 with $\delta=\frac{d}{2^{j-1}(C+1)}$, $\mathcal{F^{\prime}}:=B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$, and conditioning on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq d/2^{j-2}$ (which implies $\bar{f}\in\mathcal{F}^{\prime}$), we obtain
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}},\Upsilon_{j-1}\right)\leq 3\cdot|M_{j-1}|\exp(-n\varepsilon_{j}^{2}),$
where we define
$\varepsilon_{j}$ $=\frac{d\sqrt{\widetilde{L}(C,\sigma)}}{2^{j-1}(C+1)}.$
Further, define $a_{j}=\exp(-n\varepsilon_{j}^{2})$. Observe that $|M_{j-1}|$ is upper-bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{j-2},2(C+1))$, and by Lemma 2.9 this quantity is bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))$. The probability upper bound does not depend on $\Upsilon_{j-1}$, so we have proven that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\exp(-n\varepsilon_{j}^{2}).$
If we set $A_{j}$ to be the event that $\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}$, then we can use the property that $P(A_{J})\leq P(A_{J}|A_{J-1}^{c})+P(A_{j-1}|A_{j-2}^{c})+\dots+P(A_{3}|A_{2}^{c})+P(A_{2})$ to conclude that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\sum_{j=2}^{J}a_{j}$
$\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\frac{a_{J}}{1-a_{J}}.$
The final inequality used Lemma 2.10. Note that $a_{J}<1$ since we assume $C$ is large enough that $\widetilde{L}(C,\sigma)\neq 0$.
Note that $\frac{d}{2^{J-2}}=\frac{2\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}}.$ Now, if $n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}},2(C+1)\right)$ and $a_{J}=\exp\left(-n\varepsilon_{J}^{2}\right)<\frac{1}{2}$, which implies $(1-a_{J})^{-1}<2$, then we have
$M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}$ $=M_{\mathcal{F}}^{\operatorname{loc}}\Bigg{(}\frac{2\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}},2(C+1)\Bigg{)}\cdot\frac{a_{J}}{1-a_{J}}$
$<\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\frac{a_{J}}{1-a_{J}}$
$\leq\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\exp\left(-n\varepsilon_{J}^{2}\right)\cdot\frac{1}{1-a_{J}}$
$\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Since $a_{J}<\frac{1}{2}$ iff $n\varepsilon_{J}^{2}>\log 2$, we conclude that if

### v5

@@ Appendix B Proofs for Section 3
Appendix C Proofs for Section 4
@@
Set $A_{j}=\{\|\Upsilon_{j}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\tfrac{d}{2^{j-1}}\}$. We write $f^{\dagger}=\Upsilon_{J^{\dagger}}$ to be the output of exactly $J^{\dagger}-1$ steps, while we write $f^{\ddagger}=\Upsilon_{J^{\ddagger}+1}$ as the output of $J^{\ddagger}\geq J^{\dagger}$ steps.
If $J^{\ast}=1$, then $\varepsilon_{J^{\ast}}\asymp d$ and trivially $\mathbb{E}_{\vec{X},\vec{Y}}\|\bar{f}-f^{\ddagger}(\vec{X},\vec{Y})\|_{L_{2}(\mathbb{P}_{X})}^{2}\lesssim d^{2}$. Suppose $J^{\ast}>1$.
Part 1: Bounding $\mathbb{P}(A_{j}\cap A_{j-1}^{c})$ for $j\geq 2$.
Pick $j\geq 3$ and repeat the logic leading up to (B.7) in the proof of Theorem 3.13 to obtain
$\mathbb{P}(A_{j}\cap A_{j-1}^{c})$ $\leq\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}\mathbb{P}(\|\bar{f}-f_{i^{\ast}}\|_{L_{2}(\mathbb{P}_{X})}>(C+1)\delta),$
where $\delta=\tfrac{d}{2^{j-1}(C+1)}$. Recall $\Upsilon_{j}=f_{i^{\ast}}$ where $i^{\ast}=\mathop{\mathrm{argmin}}_{i\in[M]}\sum_{j=1}^{n}(Y_{j}-f_{i}(X_{j}))^{2}$ for $\mathcal{O}(g)=\{f_{1},\dots,f_{M}\}$ when $\Upsilon_{j-1}=g$.
Now, $\mathcal{O}(g)$ was originally formed as a (maximal) $\tfrac{d}{2^{j-1}c}$-packing of $B(g,\tfrac{d}{2^{j-2}})\cap\mathcal{F}$. We then pruned the packing but this does not affect that it is a packing. Thus, $|\mathcal{O}(g)|\leq\mathcal{M}^{\operatorname{adloc}}(g,\tfrac{d}{2^{j-2}},2c)$. But since $\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\tfrac{d}{2^{j-2}}$, we actually have $|\mathcal{O}(g)|\leq\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{d}{2^{j-3}},4c)$ by Lemma 4.2. Moreover, Lemma 3.1 states that $\mathcal{O}(g)$ forms a $\tfrac{d}{2^{j-2}c}=\tfrac{d}{2^{j-1}(C+1)}=\delta$-covering of $\mathcal{F}^{\prime}=B(g,\tfrac{d}{2^{j-2}})\cap\mathcal{F}$. Hence by Lemma 3.9,
$\mathbb{P}(A_{j}\cap A_{j-1}^{c})$
$\leq\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}3|\mathcal{O}(g)|\exp(-\Lambda(C,\sigma,L)n\delta^{2})$
$\leq\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}3\cdot\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{d}{2^{j-3}},4c)\exp(-\Lambda(C,\sigma,L)n\delta^{2}).$
Note once more that assuming $c>8$ implies $C>3$ so the application of the lemma was valid.
Next, since $\mathcal{L}(j-1)$ forms a $\tfrac{d}{2^{j-1}c}$-packing of $\mathcal{F}$ by the same lemma, it follows that the number of choices for $g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})$ is bounded by $\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{d}{2^{j-2}},2c)$. Thus,
$\mathbb{P}(A_{j}\cap A_{j-1}^{c})$
$\leq 3\cdot\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{d}{2^{j-2}},2c)\cdot\mathcal{M}^{\operatorname{adloc}}(\bar{f},\tfrac{d}{2^{j-3}},4c)\exp(-\Lambda(C,\sigma,L)n\delta^{2})$